OpenStudy (stuck-help):

METAL square root 5q^2-20=0

OpenStudy (pinklion23):

@iGreen @AnswerMyQuestions @sleepyjess @sleepyhead314

OpenStudy (pinklion23):

there maybe they can help cuz im confused lol :)

OpenStudy (stuck-help):

alright thank you

OpenStudy (pinklion23):

i have seen your question up for a while and felt bad :( you need an answer they might give it to you

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 5q^2-20=0 }\) \(\large\color{black}{ \displaystyle 5q^2-20\color{blue}{+20}=0 \color{blue}{+20}}\) \(\large\color{black}{ \displaystyle 5q^2=20 }\) \(\large\color{black}{ \displaystyle 5q^2\color{blue}{\div5}=20\color{blue}{\div5} }\) \(\large\color{black}{ \displaystyle q^2=4 }\)

OpenStudy (solomonzelman):

can you finish from here ?